### v1

# SVD-Free Low-Rank Adaptive Gradient Optimization for Large Language Models

Abstract
Low-rank optimization has emerged as a promising direction in training large language models (LLMs) to reduce the memory usage of adaptive optimizers by constraining learning to a lower-dimensional space. Prior work typically projects gradients of linear layers using approaches based on Singular Value Decomposition (SVD). However, applying SVD-based procedures individually to each layer in large models is computationally expensive and incurs additional memory costs due to storing the projection matrices. In this work, we propose a computationally efficient and conceptually simple two-step procedure to approximate SVD-based gradient projections into lower-dimensional spaces. First, we construct a complete orthogonal basis using predefined orthogonal matrices of the Discrete Cosine Transform (DCT). Second, we adaptively select basis columns based on their alignment with the gradient of each layer. Each projection matrix in our method is obtained via a single matrix multiplication followed by a lightweight sorting step to identify the most relevant basis vectors. Due to the predefined nature of the orthogonal bases, they are computed once at the start of training. During training, we store only the indices of the selected columns, avoiding the need to store full projection matrices for each layer. Our numerical experiments on both pre-training and fine-tuning tasks demonstrate the effectiveness of our dual strategy in approximating optimal low-rank projections, matching the performance of costly SVD-based methods while achieving faster runtime and reduced memory usage.
1 Introduction
The Adam optimizer (Kingma and Ba, 2014) and its regularized version AdamW (Loshchilov and Hutter, 2019) have become the standard approach for optimizing deep neural networks in various settings. With the recent increase in scale of LLMs up to billions and trillions of parameters, training with AdamW becomes more and more challenging, as its internal state requires two momentum buffers that scale with the model size. This practical problem paved the way for a line of research that focuses on reducing the memory usage of optimizer states in the context of adaptive gradient optimization. These approaches range from quantizing the states to 8 bits (Dettmers et al., 2021) to the recent GaLore optimizer (Zhao et al., 2024) inspired from the LoRA techniques (Hu et al., 2021; Lialin et al., 2023), that compresses the gradient matrix using low-rank decomposition based on SVD. Several improvements to GaLore have been proposed to enhance its performance, such as LDAdam (Robert et al., 2025), FRUGAL (Zmushko et al., 2024), FIRA (Chen et al., 2024), BAdam (Luo et al., 2024), Q-GaLore (Zhang et al., 2024b). The key aspect of GaLore and its later improvements is the low-rank decomposition based on matrix factorization, such as SVD or QR decomposition. These techniques are known to be computationally intensive as they have to be invoked for each linear layer, either at each step (for, e.g., LDAdam) or once at a few steps (for, e.g., GaLore). To mitigate the high cost of SVD, we ask: can we find an alternative projection to serve as an accurate replacement for the orthogonal matrices in SVD, that is much cheaper to compute, and portable across memory-efficient optimizers?
Contributions. In this work, we address our research question by proposing a cheaper alternative to the orthogonal matrices computed via SVD. Concretely, we propose using the orthogonal matrices from the general class of Discrete Fourier Transforms, such as the Discrete Cosine Transform (DCT), which has been successfully used in image compression for the JPEG algorithm. To the best of our knowledge, we are the first to use it in the context of low-rank adaptive gradient optimization. We summarize our contributions as follows:
  • We show the DCT matrix is an accurate replacement for the orthogonal matrices computed via SVD. We reduce the running time and improve the memory usage as we store only one DCT matrix for the entire network, computed once at the beginning of the training. Instead of storing one projection matrix for each layer, we just have to store $r$ (rank) integers, representing the indices of most significant columns for that particular layer.
  • We propose a theory-based approach to compute the most significant columns of the DCT matrix to obtain a dynamic projection matrix tailored to each layer.
  • In order to show that DCT matrices accurately approximate the orthogonal matrices of SVD, we add the DCT transformation into existing low-rank optimizers and show they recover training loss, memory usage and/or running time.
  • We propose a standalone optimizer that uses the DCT projection and optionally adds quantized error feedback for the projection error, rotates the momentum buffers such that new low-rank gradients are correctly incorporated and allows changing the low-rank subspace every step.
2 Related Work
We review the prior work in the literature focused on reducing the running time and memory usage of optimizers by compressing the gradient, especially by low-rank compression. Most approaches use individual projection matrices tailored to the gradient at each layer by invoking techniques based on quantization or matrix factorization, such as SVD, QR or PCA. One pioneering work in the context of low-rank adaptive optimization is the recent GaLore optimizer (Zhao et al., 2024) which performs SVD once at a few steps to project the gradient to a lower-dimensional space. GaLore was followed by several other optimizers that improve certain aspects of it.
The first improvement aspect is the running time, which forces GaLore to compute SVD once at a few steps in order to be feasible in practice for large models. LDAdam (Robert et al., 2025) compresses the first order momentum in AdamW, replaces SVD with a block power-iteration (Bentbib and Kanber, 2015) and performs a smooth subspace transition by rotating the first and second momentum accordingly to incorporate gradients from the same subspace at each step.
The second improvement aspect is convergence. By default, GaLore discards the projection error, which LDAdam stores and incorporates into the gradient at the next step. In contrast, the recent work FRUGAL (Zmushko et al., 2024) makes the distinction between the low-rank gradient (called state-full) and the projection error (called state-free). The state-full gradient is used in AdamW, while the state-free gradient is fed to an optimizer without state, such as SignSGD. The main idea is to leverage the remaining gradient information in the projection error since it is available at each step instead of discarding or storing it. A concurrent work to FRUGAL is FIRA (Chen et al., 2024), which aims to preserve the low-rank constraint for memory efficiency while achieving full-rank performance by proper scaling of the projection error.
In addition to improvements over GaLore, some prior work focuses on achieving memory usage similar to SGD while preserving the performance of AdamW. Apollo (Zhu et al., 2024) exploits the redundancy in the computed adaptive learning rates of AdamW and shows that element-wise learning rate is not necessary. As a consequence, they focus on computing one adaptive learning rate per channel, which significantly improves the memory usage and throughput by allowing larger batch sizes due to memory savings. They achieve this by projecting the gradient to a lower-dimensional space using a random matrix only to compute the scaling for the gradient, which remains full-rank.
Another approach worth mentioning is Online Subspace Descent Liang et al. (2024), which replaces SVD projection with Online PCA, which involves computing the projection matrix as a solution of an optimization function focused on 1) minimizing the projection error and 2) forcing the projection matrix to be orthogonal. The authors indicate that performing one step with Adam to solve this additional optimization problem is enough to obtain a qualitatively good projection matrix $P$. In contrast, our method does not introduce such overheads during training since the DCT projection is already computed at the beginning of training.
As complementary approaches to low-rank compression, there is some prior work worth mentioning due to their similar goal of reducing the memory usage and running time with a different methods than low-rank. AdamW-8bit (Dettmers et al., 2021) quantizes the states of AdamW optimizer to 8-bits using efficient CUDA kernels, with the effect of reducing the memory usage by a half compared to bfloat16 and without introducing additional computational overheads. In the same spectra of methods we find MicroAdam (Modoranu et al., 2024), where the optimizer states are compressed via sparsity and error feedback. Adam-mini (Zhang et al., 2024a) reduces optimizer states memory by assigning a single learning rate to each parameter block carefully partitioning the parameter matrix into blocks.
Despite all aforementioned approaches being similar in using SVD-based projections, they differ in a few aspects. The most important one in our view is how they handle the projection error and how often they update the low-dimensional subspace, which we clarify in Table 1.
Our approach uses DCT projection coupled with a dynamic column selection to determine a projection matrix tailored to the gradient for a particular layer, can use any update frequency and can be integrated into any optimizer, regardless of the way it handles the projection error.
Table 1: Properties of prior low-rank adaptive optimizers. The update frequency* $200$ is the default in GaLore that made the approach computationally feasible.
Low-rank Projection Type Frequency* Error
GaLore (Zhao et al., 2024) SVD 200 discard
Frugal (Zmushko et al., 2024) SVD, Random, RandPerm 200 feed to SignSGD
Fira (Chen et al., 2024) SVD 200 norm-based scaling
Subspace Descent (Liang et al., 2024) Online PCA 1 discard
LDAdam (Robert et al., 2025) Block Power Iteration 1 error feedback
DCT-AdamW (this work) DCT 1 error feedback
3 Method
This section introduces the orthogonal DCT matrix we use to replace SVD, the technique to dynamically choose columns to obtain a projection matrix tailored to the gradient at each layer, as well as the standalone optimizer DCT-AdamW we propose.
3.1 Discrete Cosine Transform
The Discrete Cosine Transform (DCT) is widely used in the signal processing and data compression (e.g., JPEG algorithm for image compression) and consists of $n$ orthogonal basis vectors whose components are cosines (Strang, 1999). There are minor variations of DCT and in this work we will use DCT-2 and DCT-3 (one is the transpose of the other). We denote the DCT-3 matrix of order $n$ by $Q\in\mathbb{R}^{n\times n}$, defined as $Q_{ij}=\sqrt{2/n}\cdot\cos\frac{i(2j+1)\pi}{2n}$, with $i,j\in[n]$, where the first row has to be divided by $\sqrt{2}$ in order for $Q$ to be orthogonal, i.e. $Q^{\top}Q=I_{n}$. We can efficiently compute $Q$ using a vectorized implementation that is fast for large values of $n$ on the GPU as follows. We first create one vector $L=[0,\cdots,n-1]^{\top}$, which is used to create the matrix $\mathcal{I}\in\mathbb{N}^{n\times n}$ by replicating $L$ on the columns of $\mathcal{I}$:
$L=\begin{bmatrix}0\\ 1\\ \vdots\\ n-1\end{bmatrix}\;\;\;\;\;\mathcal{I}=\begin{bmatrix}0&\cdots&0\\ 1&\cdots&1\\ \vdots&\cdots&\vdots\\ n-1&\cdots&n-1\\ \end{bmatrix}_{n\times n}\;\;\;\;\;Q$ $=\sqrt{\frac{2}{n}}\cos\left(\frac{\mathcal{I}\odot(2\mathcal{I}^{\top}+1)}{2n}\pi\right)$
We restate that we need to divide the first row by $\sqrt{2}$ to make $Q$ orthogonal. The computational efficiency comes from the element-wise product $\mathcal{I}\odot(2\mathcal{I}^{\top}+1)$ that actually computes the integer entries $i(2j+1)$. Once we compute $Q$, we can use it as a replacement for the SVD procedure in conjunction with the dynamic selection of columns, which we explain in detail next.
3.2 Dynamic Selection of Columns
General View. After computing the DCT matrix $Q\in\mathbb{R}^{n\times n}$, we need to compute a projection matrix $Q_{r}\in\mathbb{R}^{n\times r}$ by selecting $r$ columns from $Q$ to project the gradient $G\in\mathbb{R}^{n\times n}$ to an $r$-dimensional space $g=GQ_{r}\in\mathbb{R}^{n\times r}$. First, we compute $S=GQ$ containing the scalar products between rows of $G$ and columns of $Q$. We rank the columns of $S$ based on their $L_{1}$-norm and then pick the largest $r$ columns, which we use in $Q_{r}$. The $i^{th}$ column in matrix $S$ contains the scalar products between rows of $G$ and column $i$ of $Q$. We view these scalar products as similarities and we want to choose the columns of $Q$ with largest similarity with rows of $G$. Since the columns of $Q$ are defined as cosines of different frequencies at each component, we state that columns of $Q$ act as a filter for the rows of $G$. We choose $L_{1}$-norm because we do not want the similarities to be altered, as it would be the case for $L_{2}$-norm. Using this approach, we ensure a dynamic mechanism to choose the most appropriate columns of $Q$ for the current gradient matrix $G$. Each layer will have its own set of $r$ indices for the columns. The case illustrated so far is for a general square matrix $G$. In the context of GaLore, where $G\in\mathbb{R}^{n\times m}$, we reduce the smallest dimension of $G$ to $r$, which means we have to take into consideration two cases, based on the relationship between $n$ and $m$.
The first case is $n\geq m$, which is known as “right projection” in GaLore, where $Q\in\mathbb{R}^{m\times m}$ is the DCT-3 matrix. We compute $S=GQ$ and the corresponding $Q_{r}\in\mathbb{R}^{m\times r}$ contains the most significant $r$ columns of $S$. The second case is $n<m$, which is known as “left projection” and this time we compute $S=Q^{\top}G$, where we are interested in the rows of $Q$ and columns of $G$. Here, $Q^{\top}\in\mathbb{R}^{n\times n}$ is the DCT-2 matrix (the transpose of DCT-3, but we store only the DCT-3 matrix, $Q$) and in order to choose the columns of $Q$ we have to select $r$ indices of the largest rows in $S$.
Usually, the smallest dimension of the gradient matrix is the hidden size of the model $d_{\rm model}$. As a result, the memory overhead of replacing the SVD projection with the DCT matrix and the dynamic selection of columns is storing a single matrix $Q\in\mathbb{R}^{d_{\rm model}\times d_{\rm model}}$ for the entire model and $r$ integers for the corresponding columns from $Q$ for each layer.
3.3 DCT-AdamW
We propose DCT-AdamW, a standalone low-rank version of AdamW with DCT projection that has the option to use quantized error feedback (EF) (Seide et al., 2014; Karimireddy et al., 2019; Alistarh et al., 2018) and ensures momentum buffers integrate gradients from the same lower dimensional subspaces as in LDAdamW (Robert et al., 2025). In contrast to LDAdamW, which has to store two consecutive projection matrices per layer, we only have to store two sets of $r$ indices due to simplicity of our DCT approach coupled with the dynamic column selection technique. Prior work MicroAdam (Modoranu et al., 2024) quantized the error feedback down to 4-bits in the context of compressing the optimizer state using sparsity. In our context, we observed that the lowest resolution we can quantize EF to is 8-bits. We state our optimizer in Algorithm 1.
Algorithm 1 DCT-AdamW (right projection)
1:Input: $\beta_{1},\beta_{2},\epsilon,T,T_{u},r$
2:$m_{0},v_{0}\leftarrow 0_{n\times r},0_{n\times r}$
3:$\Xi_{1}\leftarrow 0_{n\times m}$diamond error feedback (EF) buffer
4:$Q\in\mathbb{R}^{m\times m}$diamond DCT matrix
5:$\mathcal{I}_{crt},\mathcal{I}_{prev}\leftarrow 0_{r},0_{r}$diamond column indices
6:for $t=\{1,2,\dots,T\}$ do
7:     $G_{t}\leftarrow\nabla_{\theta}f(\theta_{t})+\Xi_{t}$
8:     $R\leftarrow\textsc{UpdateSubspace}(G_{t})$
9:     $g_{t}\leftarrow G_{t}\cdot Q_{crt}$diamond projected gradient
10:     $\Xi_{t}\leftarrow G_{t}-g_{t}\cdot Q_{crt}^{\top}$diamond update EF
11:     $m_{t}\leftarrow\beta_{1}\cdot m_{t-1}\cdot R+(1-\beta_{1})g_{t}$
12:     $v_{t}\leftarrow\beta_{2}|v_{t-1}\cdot R|+(1-\beta_{2})g_{t}^{2}$
13:     $\theta_{t+1}\leftarrow\theta_{t}-\eta_{t}\frac{\hat{m_{t}}}{\epsilon+\sqrt{\hat{v_{t}}}}Q_{crt}^{\top}$
14:end for
Algorithm 2 Update Subspace procedure
1:procedure UpdateSubspace($G$)
2:     $R\leftarrow I_{r\times r}$diamond switching matrix between
3:     consecutive low-rank spaces
4:     if $t>1$ then
5:         $\mathcal{I}_{prev}\leftarrow\mathcal{I}_{crt}$
6:     end if
7:     if $(t=1)\lor(t\mod T_{u}=0)$ then
8:         re-rank the columns and
9:         $\mathcal{I}_{crt}\leftarrow\textsc{RankCols}(GQ,r)$
10:         update the switching matrix
11:         $R\leftarrow Q_{prev}^{\top}\cdot Q_{crt}$
12:     end if
13:     return $R$
14:end procedure
The sets $\mathcal{I}_{prev/crt}$ hold the indices of the $r$ columns for the previous/current projections and $Q_{prev/crt}$ contain the columns from $Q$ specified by the these two sets. $T_{u}$ represents the subspace update interval (set to 200 for GaLore and to $1$ for LDAdam). The procedure RankCols ranks the columns of $S=G_{t}\cdot Q$ as explained in Section 3.2. In order to make sure the momentum buffers integrate gradients from the same subspaces, we need to rotate $m_{t}$ and $v_{t}$ using a rotation matrix $R$ that first projects the momentum to the full-dimensional space using the previous projector and then projects it back to the current lower-dimensional subspace. We can perform this rotation directly in the $r$-dimensional space by multiplying the two projection matrices, resulting in $R\in\mathbb{R}^{r\times r}$. Rotating $v_{t}$ might introduce negative values and we apply the absolute value function to force the non-negativity of $v_{t}$. The low-rank gradient $g_{t}$ is computed by multiplying the full-rank gradient $G_{t}$ with $Q_{crt}$ and the full-rank update is obtained multipying $u_{t}$ by $Q_{crt}^{\top}$.
4 Memory Savings and Running Time
In this section we provide details about running time and memory savings after replacing the expensive SVD with DCT, coupled with the dynamic column selection approach.
4.1 Memory Savings
For simplicity, we compute the memory usage for a generic model with $L$ layers, each of size $n\times n$ for a fixed rank $r$ and provide concrete values for the Llama-2 7B model, for which $n=4096$.
SVD. To perform the low-rank decomposition for a gradient matrix $G\in\mathbb{R}^{n\times n}$, we use the SVD decomposition $G=U\Sigma V^{\top}$ and use the first $r$ columns of $U$ as a projection matrix $P=U_{:,:r}$. Storing one projection matrix $P$ for each of the $L$ layers in the neural network leads to storing additional $Lnr$ elements only for the projection matrices.
DCT. To perform low-rank decomposition using DCT, we have to store the DCT matrix $Q\in\mathbb{R}^{n\times n}$ that will be the same for all layers and $r$ additional indices at each layer, which will be used to extract the corresponding columns. Finally, the total number of elements to be stored is $n^{2}+Lr$. Usually, $n$ is the hidden dimension of the model $d_{\rm model}$, which the memory usage directly depends on.
Practical Memory Usages. We compute the memory usage for a Llama-2 7B model to understand the practical implications of using a constant orthogonal matrix compared to SVD-based projections. Table 2 shows the memory usage of our approach is not sensitive to rank, since we only store $r$ integers instead of individual projection matrices with $r$ columns. Instead, the memory usage of SVD depends on both hidden dimension of the model $n=d_{\rm model}$ and the rank, since $r$ columns have to be stored. This translates to memory savings for our approach for LLMs. In DDP settings, we store one projection matrix for each GPU. However, when the model size becomes too large, GPUs can store different columns of the matrix to prevent materializing the full projection matrix on each GPU.
Table 2: Practical memory usage for low-rank projection matrices for Llama-2 7B.
SVD DCT
$r=32$ $r=256$ $r=512$ $r=32$ $r=256$ $r=512$
bfloat16 56 MiB 448 MiB 896 MiB 32.03 MiB 32.22 MiB 32.44 MiB
float32 112 MiB 896 MiB 1792 MiB 64.03 MiB 64.22 MiB 64.44 MiB
4.2 Running Time
Once the projection matrices $P$ are computed and stored, the computational complexity of the up/down-projection operations is the same as for SVD, as each involves a matrix multiplication between the projection matrix (or its transpose) and the full/low-rank gradient. However, since our projection matrices are based on the Discrete Cosine Transform (DCT), we can leverage fast Fourier transform (FFT) algorithms to reduce the complexity of the up/down-projection operations from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n\log n)$. It is important to note that our method for computing the projection matrices is fundamentally different from existing approaches. In this work, we primarily focus on comparing the runtime of our method with that of SVD-based approaches.
For a generic matrix $G\in\mathbb{R}^{n\times n}$, the computational complexity of SVD is $\mathcal{O}(n^{3})$. This complexity has to be incurred for each layer once at $T_{u}$ steps because this procedure is too expensive to be run with higher frequency. On the other hand, our approach computes $S=GQ$, ranks the columns in $S$ by computing $L_{1}$-norm per column and selects the top-$r$ columns to form the projection matrix $P$ to be used for up/down-projection. At steps where we recompute the projection, we can avoid an additional matrix multiplication between $P$ and the full rank gradient matrix $G$ by returning the largest $r$ columns of $S$ as the low-rank gradient. Overall, our approach has a lower computational overhead compared to the SVD approach, which can also be seen in the running time row of results in experimental section.
5 Theoretical Guarantees
In this section, we provide a theoretical justification for our two-step procedure to approximate SVD-based gradient projections. First, we rigorously show that adaptively selecting columns of an orthogonal matrix based on their alignment with the gradient matrix is an optimal strategy for minimizing the reconstruction error. This approach leads to a contractive compression scheme, which is commonly exploited in the analysis of compressed adaptive optimization algorithms. Next, to justify the specific use of the DCT matrix, we demonstrate that it naturally serves as a linear approximation of the left or right eigenbases of the gradient matrices.
5.1 Optimality of norm-based ranking procedure
Let $G\in\mathbb{R}^{n\times m}$ be the gradient matrix and $Q\in\mathbb{R}^{n\times n}$ is orthogonal, namely $QQ^{\top}=I_{n}$ (without loss of generality, we consider left multiplication). For a given rank $r\leq n$, let $Q_{r}$ be a $n\times r$ sub-matrix of $Q$ composed of $r$ columns. We are interested in the reconstruction error between $G$ and $Q_{r}Q_{r}^{\top}G$, where $Q_{r}^{\top}$ performs projection and $Q_{r}$ performs back-projection. While we may choose any matrix norm to measure the reconstruction error, the standard Frobenius norm makes the derivation cleaner. Specifically, using the definition of the Frobenius norm $\|G\|_{\rm F}^{2}={\rm tr}(G^{\top}G)$, linearity of trace and $Q_{r}^{\top}Q_{r}=I_{r}$ due to orthogonality of $Q$, we decompose the reconstruction error:
$\|G-Q_{r}Q_{r}^{\top}G\|_{\rm F}^{2}$ $=$ ${\rm tr}(G^{\top}(I-Q_{r}Q_{r}^{\top})^{\top}(I-Q_{r}Q_{r}^{\top})G)={\rm tr}(G^{\top}(I-Q_{r}Q_{r}^{\top})G)$
$=$ ${\rm tr}(G^{\top}G-G^{\top}Q_{r}Q_{r}^{\top}G)={\rm tr}(G^{\top}G)-{\rm tr}(G^{\top}Q_{r}Q_{r}^{\top}G)$
$=$ $\textstyle\|G\|_{\rm F}^{2}-\|Q_{r}^{\top}G\|_{\rm F}^{2}=\|G\|_{\rm F}^{2}-\sum_{i=1}^{r}\|q_{i}^{\top}G\|_{2}^{2},$
where $q_{i}$’s are the columns of $Q_{r}$. From this identity, we conclude that to minimize the reconstruction error, the optimal strategy is to maximize the alignments $\|q_{i}^{\top}G\|_{2}^{2}$ for all selected columns $q_{i}$. Furthermore, since $\|q_{1}^{\top}G\|_{2}^{2}+\|q_{2}^{\top}G\|_{2}^{2}+\dots+\|q_{n}^{\top}G\|_{2}^{2}=\|Q^{\top}G\|_{\rm F}^{2}={\rm tr}(G^{\top}QQ^{\top}G)={\rm tr}(G^{\top}G)=\|G\|_{\rm F}^{2},$ following the optimal strategy of selecting $r$ columns from $Q$, we have $\|G-Q_{r}Q_{r}^{\top}G\|_{\rm F}^{2}=\|G\|_{\rm F}^{2}-\sum_{i=1}^{r}\|q_{i}^{\top}G\|_{2}^{2}\leq\|G\|_{\rm F}^{2}-\frac{r}{n}\sum_{i=1}^{n}\|q_{i}^{\top}G\|_{2}^{2}=\left(1-\frac{r}{n}\right)\|G\|_{\rm F}^{2}.$ Thus, the proposed norm-based selection strategy for any orthogonal matrix $Q$ induces a low-rank compression scheme that is contractive with a factor $1-\nicefrac{{r}}{{n}}$. Contractivness of the compression scheme is the key property in the convergence analysis of compressed optimization (Stich et al., 2018; Richtárik et al., 2021; Li et al., 2022; Modoranu et al., 2024; Robert et al., 2025).
We can extend the above argument for any $p$-norm over vectorized matrices, for which the Frobenius norm is the special case of $p=2$. Then, we have
$\|G-Q_{r}Q_{r}^{\top}G\|_{p}$ $=\textstyle\left\|\sum_{i=1}^{n}q_{i}q_{i}^{\top}G-\sum_{i=1}^{r}q_{i}q_{i}^{\top}G\right\|_{p}=\left\|\sum_{i=r+1}^{n}q_{i}q_{i}^{\top}G\right\|_{p}\overset{(a)}{\leq}\sum_{i=r+1}^{n}\left\|q_{i}q_{i}^{\top}G\right\|_{p}$
=$\overset{(b)}{=}\textstyle\sum_{i=r+1}^{n}\left\|q_{i}\right\|_{p}\left\|q_{i}^{\top}G\right\|_{p}\overset{(c)}{\leq}\max(1,n^{\frac{1}{p}-\frac{1}{2}})\sum_{i=r+1}^{n}\left\|q_{i}^{\top}G\right\|_{p},$
where (a) follows from the triangle inequality of the $p$-norm, (b) follows from the definition of $p$-norms for matrices, namely $\|uv^{\top}\|_{p}=(\sum_{i,j}u_{i}^{p}v_{j}^{p})^{\nicefrac{{1}}{{p}}}=(\sum_{i}u_{i}^{p}\sum_{j}v_{j}^{p})^{\nicefrac{{1}}{{p}}}=\|u\|_{p}\|v\|_{p}$ for two column-vectors $u$ and $v$ of the same size, (c) follows from the relationship between $\ell_{p}$ norms and that $\|q_{i}\|_{2}=1$, i.e., $\|v\|_{p}\leq n^{\frac{1}{p}-\frac{1}{2}}\|v\|_{2}$ if $p\leq 2$ and $\|v\|_{p}\leq\|v\|_{2}$ if $p\geq 2$.
5.2 DCT as linear approximation of the gradient eigenbasis
Given a real-valued gradient matrix $G\in\mathbb{R}^{n\times m}$ and its SVD decomposition $G=U\Sigma V^{\top}$, our goal is to find fast approximation of (without loss of generality) its left eigenvectors stacked in $U\in\mathbb{R}^{n\times n}$. As $GG^{\top}=U\Sigma^{2}U^{\top}$, it is equivalent to approximate the eigenvectors of symmetric matrix $GG^{\top}$.
Our argument starts from a linear algebra decomposition result, originally motivated by the optical information processing literature to factorize linear transformations that can be implemented optically (Müller-Quade et al., 1998; Schmid et al., 2000). The result states that any square matrix $M$ of shape $n\times n$ over the complex numbers $\mathbb{C}$ can be decomposed into a product of diagonal and circulant matrices, i.e., $M=D_{1}C_{2}D_{3}\dots D_{2k-3}C_{2k-2}D_{2k-1}$, where $D$’s are diagonal matrices and $C$’s are circulant matrices. Circulant matrices are a special class of Toeplitz matrices in which each row is a cyclic right shift of the previous one as shown below:
$C=\begin{bmatrix}c_{0}&c_{1}&c_{2}&\cdots&c_{n-1}\\ c_{n-1}&c_{0}&c_{1}&\cdots&c_{n-2}\\ \vdots&\vdots&\cdots&\vdots&\vdots\\ c_{1}&c_{2}&\cdots&c_{n-1}&c_{0}\\ \end{bmatrix},\quad F=\frac{1}{\sqrt{n}}\left[w^{ij}\right]_{i,j=0}^{n-1},\text{ where }w=e^{\frac{2\pi i}{n}}.$ (1)
[…]
$GG^{\top}=(FD_{1}F^{*})D_{2}(FD_{3}F^{*})D_{4}\cdots D_{2k-2}(FD_{2k-1}F^{*}).$ (2)
Notice the analogy between this matrix decomposition and the Taylor expansion for functions. Since the space of matrices is finite-dimensional, the signal-matrix can be recovered using finitely many “variable” $D$’s, in contrast to the infinite series required for function expansions. Keeping this analogy in mind, just as loss functions are linearly approximated at each iteration in first-order optimization algorithms, we consider a “linear” approximation of the decomposition (2) by including only a single variable $D$, i.e., $GG^{\top}\approx FD_{1}F^{*}$.
This directly implies that we approximate the eigenvectors $U$ by the DFT matrix $F$. However, since the matrix $GG^{\top}$ is real and symmetric by design, its eigenvalues are real, and the real part $\text{Re}(F)$ also forms an approximate eigenbasis that better aligns with $U$. Finally, we observe that the real part $\text{Re}(F)$ corresponds to the discrete cosine transform (DCT), up to minor variations.
6 Experiments
We now present our numerical results. Our main goal is to show that DCT projection can recover the performance of SVD-based projection and have lower memory footprint and running time.
For pre-training, we use a Llama model (Touvron et al., 2023) with 800 million parameters and train it on the C4 dataset using 20/100 tokens/parameter, equivalent to 16/80B tokens. We replace the SVD projection in Frugal and Fira with DCT matrix to directly compare the optimizer behavior when the only difference is the projection. We use a learning rate tuned for the full-rank AdamW baseline to avoid the expensive learning rate tuning process for each individual projection. As a result, in certain cases, the DCT projection would require further learning rate tuning. In all settings, our main comparison is against the SVD-like projection and we will also add the full-rank AdamW results for reference. For fine-finetuning, we use Llama-2 7B and report the accuracy on GSM-8k dataset (Cobbe et al., 2021) using Frugal and Fira with SVD and DCT projections, as well as on LDAdamW and our DCT-AdamW implementation.
6.1 Pre-training (PT)
We train the Llama-800M model from scratch using 16B and 80B tokens, equivalent to 20 and 100 tokens per parameter, from C4 dataset (Raffel et al., 2020). The hidden dimension of the model is $2048$ and we use rank $r=256$ for all runs, as well as learning rate $\eta=7.5e-5$, which was chosen for the full-rank AdamW, proportionally to the inverse square root of the number of parameters and kept fixed for all runs.
PT with FRUGAL. We integrate the DCT matrix into the FRUGAL optimizer and compare against the SVD, RandPerm and Random projections. RandPerm uses a random permutation as a projection matrix, while Random generates a random, semi-orthogonal matrix. In Figure 1(a) we show the training loss for all these runs. In the zoomed-in plot for the last 5000 training steps we see that SVD projection recovers the performance full-rank AdamW, while the DCT projection is a good approximation of the SVD, which supports our claim. We present our numerical results in Table 3. We would like to emphasize the running time reduced by 1h 48m ($\approx 22.6\%$) compared to the SVD projection, as well as the memory usage reduced by about 2.2GB ($\approx 3.5\%$), while the increase in perplexity is less than one point. In comparison to RandPerm and Random projections, the runtime is on par, while train and validation perplexities are lower by approximately one point in the favor of DCT, illustrating again the benefit of our approach.
PT with FIRA. We integrate DCT into FIRA optimizer and compare against SVD. In Figure 1(b) we show the training loss, where our focus is on the comparison between SVD and DCT projections. We observe that DCT consistently yields lower training loss compared to the SVD projection, which is also visible in the numerical results in Table 3, translated to lower perplexities in the favor of DCT. Moreover, the memory usage is smaller by 2GB ($\approx 3\%$) and the running time is lower by 1h 54m ($\approx 23.8\%$).
Table 3: Pre-training Results for AdamW, FRUGAL and FIRA with different projections using 20 tokens/parameter. DCT is a good approximation to SVD with lower runtime and memory. AdamW is the full-rank optimizer and is added for reference.
FRUGAL FIRA
AdamW SVD DCT RandPerm Random SVD DCT
Train PPL 15.55 15.35 15.63 16.52 17.02 19.37 18.88
Val. PPL 14.05 14.02 14.23 15.18 15.61 17.67 17.30
Mem. (GiB) 73.49 65.70 63.50 65.44 63.72 68.44 66.48
Time 7h 54m 9h 45m 7h 57m 7h 56m 7h 56m 9h 53m 7h 59m
Figure 1: Pre-training Llama-800M using FRUGAL and FIRA on 16B tokens from C4.
[…]
Table 4: Pre-training results on Llama-800M for AdamW, LDAdamW and DCT-AdamW with 100 tokens/parameter. AdamW is the full-rank optimizer and is added for reference.
AdamW LDAdamW DCT-AdamW
Train PPL 12.88 15.10 14.95
Val. PPL 11.73 13.91 13.69
Mem. (GiB) 73.72 72.10 57.82
Time 1d 13h 22m 2d 1h 24m 1d 15h 17m
Figure 2: Pre-training Llama-800M with DCT-AdamW and LDAdam on 80B tokens from C4.
6.2 Fine-tuning (FT)
We finetune Llama-2 7B on GSM-8k dataset using SVD and DCT projections for FRUGAL and FIRA optimizers, as well as LD-AdamW and DCT-AdamW. We show our results for ranks $r\in\{32,512\}$ in Table 5.
FT with FRUGAL. Both projections yield similar training loss for both rank values, which is an indication that DCT is indeed a good approximation for SVD. Despite having different training loss, the SVD projection achieves the same accuracy for both ranks. It is surprising that DCT yields better accuracy for lower rank compared to higher rank. In terms of memory usage, as expected for this large model, the DCT projection saves 8GB of memory ($\approx 23.4\%$) for $r=32$, while for $r=512$ the reduction in memory is only 6.3GB ($\approx 18.2\%$). The running time is reduced by roughly 35m ($\approx 75\%$) for both ranks.
Table 5: Fine-tuning results for Llama-2 7B on the GSM-8K dataset.
FRUGAL FIRA LD/DCT-AdamW
rank 32 rank 512 rank 32 rank 512 rank 32 rank 512
SVD DCT SVD DCT SVD DCT SVD DCT LD DCT LD DCT
Train Loss 0.046 0.059 0.051 0.071 0.123 0.209 0.094 0.192 0.261 0.176 0.101 0.208
Acc. (%) 33.81 35.93 33.81 34.26 32.15 32.45 34.27 35.25 31.38 30.09 35.61 35.17
Mem. (GB) 39.61 31.59 40.68 34.41 32.72 31.29 35.41 34.11 32.08 31.17 35.58 34.35
Time 1h 22m 46m 1h 19m 47m 1h 8m 1h 1h 9m 1h 55m 48m 1h 8m 48m
FT with FIRA. The DCT projection yields larger training loss compared to SVD. DCT recovers the accuracy, outperforms the SVD for large rank, and on average reduces the memory usage by $\approx 1.35$ GB and the running by $\approx$10m.
FT with LDAdamW & DCT-AdamW. In this setting we compare the DCT projection with the block power-iteration used in LDAdamW as an approximation to SVD, both without EF. The training loss achieved is smaller for DCT on $r=32$ and larger for $r=512$ compared to LDAdamW. LDAdamW achieves more than 1% accuracy for $r=32$ and comparable accuracy for $r=512$. Regarding memory, DCT-AdamW uses $\approx$1 GB less memory, while achieving a speedup of about 20m for $r=512$. We would like to emphasize that EF does not help DCT-AdamW in comparison to LDAdamW for $r=32$, the accuracy of DCT-AdamW with EF being 29.11% vs 32.53% for LDAdam with EF for $r=32$, while for $r=512$ the accuracy can be recovered, scoring 35.33% compared to 35.86% for LDAdamW with EF.
7 Limitations
While our approach shares the underlying principle of low-rank gradient approximation with optimizers like GaLore, FRUGAL, FIRA, and LDAdamW, it is distinctly designed with the computational challenges and scalability demands of large language models (LLMs) in mind. As such, our evaluation and optimization strategy are tailored to this setting, and we leave exploration in domains like computer vision to future work.
Appendix A Hyper-parameters and Reproducibility
In this section we provide details about the hyper-parameters we used to obtain the results in the main paper, including hardware description, in order to reproduce the results.
A.1 Pre-training
A.1.1 Model
We use the llm-baselines [^2] https://github.com/epfml/llm-baselines, a public repository that implements the Llama architecture from scratch and allows creating models with different number of parameters. We use the version with 822M parameters. We would like to emphasize that the number of parameters stated here does not include the embeddings and classification head, which have both size 65.5M (if included, then total model size is 953M parameters). The model has 16 Llama blocks with embedding size 2048 and 16 attention heads.
For our experiments, we train the models from scratch using 20 and 100 tokens per parameter. The linear layers are trained using the optimizers mentioned in the paper, while the embeddings, classification head and layer normalization layers are trained with AdamW.
A.1.2 Hyper-Parameters
Due to the lengthy training time of pretraining, we performed tuning over a limited range of learning rates for all optimizers. We selected learning rates that scale with $\propto 1/\sqrt{d}$, where $d$ represents the model size. Our learning rate grid consisted of $\{7.5e-5,1e-4,2.5e-4\}$ for the experiments involving 20 tokens per parameter. For settings with 100 tokens per parameter, we refrained from performing any grid search because the process would be computationally intensive, and consequently, we maintained the learning rate fixed at $57.5e-$ 5.
A.2 Fine-tuning
A.2.1 Model
We use the llm-foundry [^3] https://github.com/mosaicml/llm-foundry repository to finetune Llama-2-7B [^4] https://huggingface.co/meta-llama/Llama-2-7b and Llama-3 8B [^5] https://huggingface.co/meta-llama/Meta-Llama-3-8B pretrained by Meta and we download the models from HuggingFace.
A.3 Hardware
We use NVidia H100 80GB GPUs for all experiments. For pre-training, we used PyTorch DPP with 8 GPUs and for fine-tuning we used only one GPU per experiment.
Appendix B Extended Experiments
B.1 Pre-training
We extend the experimental section in the main paper by adding experiments with Apollo optimizer for pre-training in the setting with 20 tokens per parameter. We would like to emphasize that FRUGAL uses the same learning rate as in the main paper (i.e. $57.5e-$ 5), while the other optimizers (FIRA, Apollo and GaLore) have slightly larger learning rate obtained via grid search (i.e. $42.5e-$ 4, compared to $57.5e-$ 5 in the main paper). In these experiments we use rank $r=256$.
Table 6: Extended pre-training experiments for AdamW, FRUGAL and FIRA with different projections using 20 tokens/parameter. DCT is a good approximation to SVD with lower runtime and memory. LR stands for Learning Rate.
Optimizer Projection LR Train PPL Val. PPL Mem. (GiB) Time
FRUGAL SVD $57.5e-$ 5 15.35 14.02 65.70 9h 45m
DCT 15.63 14.23 63.50 7h 57m
RandPerm 16.52 15.18 65.44 7h 56m
Random 17.02 15.61 63.72 7h 56m
FIRA SVD $42.5e-$ 4 19.37 17.67 68.44 9h 53m
DCT 18.88 17.30 66.48 7h 59m
Apollo Random $42.5e-$ 4 17.00 15.53 67.40 7h 56m
GaLore SVD $42.5e-$ 4 17.82 16.32 69.90 9h 32m
DCT-AdamW DCT $42.5e-$ 4 18.22 16.66 68.66 7h 42m
Figure 3: Pre-training Llama-800M with FRUGAL (SVD, DCT, RandPerm, Random), FIRA (SVD, DCT), Apollo (Random) and GaLore (SVD).
B.2 Fine-tuning
We extend the experimental section in the main paper by adding experiments on Llama-2-7b on VIGGO and SQL datasets and on Llama-3-8b on GSM-8k, VIGGO and SQL. We train for 3 epochs on GSM-8k and VIGGO and for one epoch on SQL.
B.2.1 Llama-2-7b on Viggo & SQL
Table 7: Fine-tuning results for Llama-2 7B on the VIGGO dataset. (*) The score $a/b$ for LD means that $a$ didn’t have EF and $b$ had EF. (**) The score $a/b/c$ for DCT means that $a$ was run without rotating states and $T_{u}=200$, $b$ was run with rotating states and $T_{u}=200$ and $c$ was run with rotating states and $T_{u}=1$, all without EF. Red values mean we would need an extended learning rate grid because the reported accuracy was obtained at the end of the grid.
Method Rank Projection Acc. (%) Mem. (GB) Time
FRUGAL 32 SVD 95.10 39.69 50m
DCT 95.09 31.49 32m
512 SVD 89.21 40.67 48m
DCT 96.08 34.11 33m
FIRA 32 SVD 87.11 32.62 48m
DCT 91.87 31.38 35m
512 SVD 92.15 35.16 50m
DCT 94.96 34.12 33m
GaLore 32 SVD 81.65 32.38 47m
DCT 41.45 31.38 34m
512 SVD 94.54 35.18 48m
DCT 80.39 34.13 32m
LD/DCT-AdamW 32 LD* 67.92 / 87.39 31.89 40m
DCT** 44.25 / 75.07 / 85.15 31.15 35m
512 LD* 86.97 / 96.91 35.11 48m
DCT** 43.13 / 16.67 / 42.85 34.16 32m
Table 8: Fine-tuning results for Llama-2 7B on the SQL dataset
Method Rank Projection Train Loss Acc. (%) Mem. (GB) Time
FRUGAL 32 SVD 0.049 22 39.47 1h 28m
DCT 0.035 11.8 31.51 1h 6m
512 SVD 0.060 46.4 40.77 1h 28m
DCT 0.021 29.9 34.28 1h 3m
FIRA 32 SVD 0.044 81.1 32.39 1h 28m
DCT 0.062 49.3 31.16 1h 3m
512 SVD 0.043 57.9 35.29 1h 29m
DCT 0.026 87.9 34.05 1h 3m
GaLore 32 SVD 0.044 80.7 32.17 1h 26m
DCT 0.039 81.6 31.16 1h 1m
512 SVD 0.068 78.5 35.30 1h 26m
DCT 0.037 86 34.06 1h 1m
LD/DCT-AdamW 32 LD 0.053 / 0.094 42.2 / 66 31.74 1h 8m
DCT 0.038 / 0.03 / 0.036 77.1 / 79.6 / 56.4 31.05 1h 4m
512 LD 0.04 / 0.032 53.9 / 32.7 35.1 1h 25m
DCT 0.032 / 0.035 / 0.046 86.9 / 36.3 / 87.9 34.09 1h 1m
B.2.2 Llama-3-8b on GSM-8k, Viggo & SQL
Table 9: Fine-tuning results for Llama-3 8B on the GSM-8K dataset. (*) means the best performance was obtained on the largest learning rate in the grid (larger learning rate might help). See also Appendix B.2.3.
Method Rank Projection Train Loss Acc. (%) Mem. (GB) Time
FRUGAL 32 SVD 0.066 35.70 52.25 1h 6m
DCT 0.049 39.04 42.35 51m
512 SVD 0.077 38.74 54.35 1h 8m
DCT 0.054 40.18 44.74 50m
FIRA 32 SVD 0.233 59.74 41.69 1h 8m
DCT 0.237 59.97 41.38 50m
512 SVD 0.192 59.97 45.91 1h 8m
DCT* 0.122 52.61 44.53 53m
GaLore 32 SVD 0.299 61.26 41.70 1h 4m
DCT* 0.334 55.5 41.38 48m
512 SVD 0.182 60.65 45.87 1h 4m
DCT* 0.317 56.02 44.51 48m
LD/DCT-AdamW 32 LD 0.262 59.97 40.73 54m
DCT 0.220 / 0.239 / 0.204 53.6 / 56.10 / 47.46 41.34 49m
512 LD* 0.115 56.78 45.12 1h 5m
DCT* 0.240 / 0.282 / 0.126 57.46 / 52.84 / 42.15 44.51 50m
Table 10: Fine-tuning results for Llama-3 8B on the VIGGO dataset. See also Appendix B.2.3.
Method Rank Projection Acc. (%) Mem. (GB) Time
FRUGAL 32 SVD 94.67 52.20 45m
DCT 93.98 42.3 34m
512 SVD 91.6 54.3 46m
DCT 93.84 44.42 34m
FIRA 32 SVD 95.94 41.64 45m
DCT 96.5 41.32 34m
512 SVD 95.94 45.85 45m
DCT 95.08 44.45 34m
GaLore 32 SVD 95.23 41.65 45m
DCT 93.98 41.32 33m
512 SVD 95.52 45.81 45m
DCT 94.81 44.41 33m
LD/DCT-AdamW 32 LD 78.99 40.66 40m
DCT 95.24 41.3 35m
512 LD 84.87 45.06 48m
DCT 95.65 44.45 36m
Table 11: Fine-tuning results for Llama-3 8B on the SQL dataset. See also Appendix B.2.3.
Method Rank Projection Train Loss Acc. (%) Mem. (GB) Time
FRUGAL 32 SVD 0.056 77.4 51.23 1h 26m
DCT 0.071 76.9 42.30 1h 8m
512 SVD 0.084 73.5 54.32 1h 28m
DCT 0.073 34.6 44.67 1h 10m
FIRA 32 SVD 0.062 78.4 41.64 1h 26m
DCT 0.054 79.5 41.32 1h 10m
512 SVD 0.054 79.2 45.86 1h 27m
DCT 0.048 79.5 44.5 1h 11m
GaLore 32 SVD 0.066 77.7 41.64 1h 22m
DCT 0.071 77.2 41.13 1h 5m
512 SVD 0.063 79.3 45.85 1h 23m
DCT 0.058 79.5 44.29 1h 6m
LD/DCT-AdamW 32 LD 0.032 79.7 40.47 1h 10m
DCT 0.074 / 0.041 / 0.063 64.4 / 77.3 / 78.1 41.3 1h 9m
512 LD 0.073 80.5 45.09 1h 30m
DCT 0.047 / 0.056 / 0.071 78.8 / 77.1 / 77.9 43.52 1h 8m
B.2.3 Particularity of Llama-3-8B architecture
We would like to mention a significant different in the architecture of Llama-3-8B in comparison with Llama-2-7B that disadvantages the DCT projection.
First of all, the main idea of our work is to store one DCT matrix of shape $n\times n$, where $n$ is the embedding dimension. For Llama-2-7b, we can easily store a single DCT matrix because the smallest dimension of all linear layers is $n=4096$. However, in the case of Llama-3-8B, the MLP projections for the values (i.e. v_proj) have shapes $(1024,4096)$, where 4096 is the embedding dimension. According to the GaLore rules, these layers should be projected on the left because the smallest dimension is 1024. Performing a left projection would mean we need a DCT matrix of order $1024$ such that the dimension of the projected gradient for this layer is $(r,4096)$. We do not want to store a second DCT matrix of shape $(1024,1024)$ and as a consequence we decided to perform right projection in this case, since we already have the DCT matrix of shape $(4096,4096)$ already allocated. As a result, the projected layer will have the shape $(1024,r)$. We would like to emphasize that this design choice affects the results for Llama-3-8B because the MLP layers v_proj are embedded to a much lower dimensional space and has the potential of becoming a bottleneck in the network.
The idea we suggest to solving this issue without allocating a new DCT matrix is to use the first 1024 rows and 1024 columns of the DCT matrix of size $(4096,4096)$ (properly scaled with a factor that is proportional to $\sqrt{4096/1024}$) such that the resulting submatrix is orthogonal and perform the left projection as in the original GaLore.